Nicholas Thompson begins a long-distance walk at a rate of 4 mph. Forty-five minutes after he leaves, his wife, Kristin, realizes that Nicholas has forgotten his wallet. Kristin gets on her bicycle and begins riding at a rate of 24 mph along the same path that Nicholas took.
Calculate how long it will take Kristin to catch Nicholas.
Calculate how far from their home Kristin will catch Nicholas.

Answer by mananth(16949)

Nicholas's speed = 4mph
distance = speed * time
in 3/4 hour he travels 4*3/4 miles = 3 miles
...
wife's speed = 24 mph.
every hour she catches up 24-4 = 20 miles.
to catch up 3 miles she will take 3/20 hours
=3/20 * 60 minutes
= 9 minutes